# combined_file.bib Non-classical Temperature Dependence of Chirality-Induced Magnetization and Its Implications for RNA’s Homochirality

### v1

@@ Abstract
The single-handedness of biomolecules is a characteristic feature of life on Earth, yet achieving and maintaining a homochiral prebiotic network is a significant challenge. In our previous studies, we reached homochirality in an RNA precursor, ribo-aminooxazoline (RAO), using magnetized magnetite surfaces as chiral reagents, due to the chiral-induced spin selectivity (CISS) effect. We further demonstrated that RAO can induce net magnetization on previously demagnetized surfaces, indicating a self-reinforcing feedback between chiral molecules and magnetic surfaces. However, these processes depend on spin interactions that generally weaken at higher temperatures, raising concerns about the robustness of our mechanisms in natural settings. Additionally, the temperature dependence of CISS remains poorly understood. To address these questions, we investigated the temperature dependence of chirality-induced magnetization. Contrary to classical expectations, we observed a significant increase in induced net magnetization with increasing temperatures, suggesting a phonon-assisted process. Concurrently, we noted a corresponding increase in the relative yield of RAO during its prebiotic synthesis. Our results support a vibronic contribution to CISS and indicate that spin-controlled processes leading to RNA homochirality can occur reliably and effectively over a range of temperatures likely present in prebiotic environments.
Life is intrinsically linked to planetary environments, with prebiotic chemistry constrained by the environmental conditions [sasselov2020origin]. These planetary environments present a wide range of conditions that vary over different timescales, from diurnal fluctuations to gradual geological shifts. Consequently, the processes leading to the formation of the first cells must be both robust and adaptable to these dynamic conditions. A critical process in life’s origins is the emergence of homochirality, the single-handedness of biomolecules. Establishing homochirality at the monomer level in prebiotic synthesis is likely essential for the emergence of a functional genetic polymer, like RNA, with high yields and selectivity [joyce1984chiral, ozturk2024new]. Nevertheless, despite its significance, achieving and maintaining a homochiral prebiotic network has long been an unresolved problem for prebiotic chemistry [blackmond2019origin]. Numerous hypotheses have been proposed for chiral symmetry breaking and the amplification of enantiomeric excess [sallembien2022possible]. Yet, none of these have demonstrated a robust mechanism for breaking chiral symmetry in a prebiotic molecule and amplifying it to homochirality from fully racemic starting materials, whereby the chiral information can be effectively propagated to establish a homochiral network (Figure 1b).
[…]
@@
Central to our findings is the chiral-induced spin selectivity (CISS) effect—an empirical phenomenon revealing a strong interaction between electron spin and molecular chirality, even at room temperature [naaman2012chiral, naaman2019chiral, bloom2024chiral]. The strong spin-chirality coupling due to CISS [ziv2019afm, safari2024enantioselective] allows for achiral magnetic minerals to function as chiral reagents for chemical reactions, which can be controlled by electron spin [banerjee2018separation, bloom2020asymmetric, metzger2021dynamic]. Although various independent measurements confirm CISS-related effects, a complete microscopic understanding of CISS remains an ongoing challenge [evers2022theory]. Notably, recent theoretical work by Fransson emphasizes the role of electron correlations in explaining the high levels of spin-polarization observed in experiments [fransson2019chirality].
@@
Additionally, the effect of temperature on CISS through electron-phonon interactions presents an intriguing complexity [bloom2024chiral, kim2023chiral, fransson2020vibrational], with experimental results showing trends that appear incompatible with each other [das2022temperature, qian2022chiral, alwan2023temperature, yang2023real]. This complexity is partly due to the fact that temperature also affects the material and molecular properties involved, complicating efforts to isolate the CISS component of the observed effects. Moreover, as our previous findings indicate that spin-chirality interactions may play a role in life’s homochirality in natural settings, we must consider temperature as an environmental variable given its significant variability in natural environments (Figure 1d). To address these challenges, we examined the temperature dependence of chirality-induced avalanche magnetization in chiral RAO crystals on magnetic surfaces stable within our temperature range, as well as the effect of temperature on the prebiotic synthesis of RAO.
@@
Contrary to classical intuition about spin-controlled processes (Figure 1d), our findings demonstrate that the net magnetization induced by chiral molecules due to CISS, through their spin-specific interactions with the magnetic surface, significantly increases with rising temperatures (Figure 2 and Figure 3). This supports a vibronic contribution to CISS, indicating that phonon-assisted couplings play a significant role in the observed spin polarizations due to the CISS effect [fransson2020vibrational]. These findings shed light on the temperature dependence of CISS, which remains an active area of research with seemingly incompatible findings. Commensurable with these, we observed a corresponding increase in the relative yield of RAO during its prebiotic synthesis as temperature rises (Figure 4). Our results, therefore, reveal the effect of an essential component of CISS, temperature, with important implications for life’s homochirality—suggesting that spin-controlled processes contributing to RNA’s homochirality via CISS can robustly occur across a range of conditions likely present in natural environments.
@@
How can we account for the increased effectiveness of CISS with temperature without involving microscopic details? We will start with a general thermodynamic picture and incorporate microscopic analyses in the discussions.
CISS relates to the spin-filtering of electrons, meaning that chiral molecules reduce entropy by decreasing the number of available microstates of the electronic spin distribution in the microcanonical ensemble. In this framework, temperature is a derived quantity that is inversely proportional to the change in entropy. The system of interest can be considered as the electronic spin distribution of the surface molecules and the interface. According to the first law of thermodynamics, the internal energy of an isolated system remains constant. Therefore, any work done by the system must be supplied as heat entering the system. Chiral molecules perform work by filtering spin, and we can think of them as chiral molecular Carnot engines or “chiral engines”, as illustrated in (Figure 1c). To balance the work performed, these systems must absorb heat from their surroundings. Thus, chiral engines operate more effectively when heat is provided.
@@
This thermodynamic picture is an intuitive way of visualizing the temperature dependence of CISS. However, the effect itself is a complex phenomenon involving spin-orbit coupling, electron correlations, and electron spin-phonon interactions. Fransson investigated the role of phonons and vibronic interactions in CISS, constructing a microscopic theory of CISS beyond the Born–Oppenheimer approximation—elucidating the exchange of angular momentum as electrons traverse a helical trajectory [fransson2020vibrational, fransson2022charge]. This model is introduced in our discussions and will be elaborated upon in a forthcoming work by Fransson et al. Our experimental results align with this microscopic framework, indicating that phonon-chirality couplings, amplified by temperature, play an essential role in CISS.
Results
Temperature dependence of chirality-induced magnetization
We studied the effect of temperature on chirality-induced magnetization using a magneto-optical Kerr effect (MOKE) microscope and a Hall device. Both measurements showed an increase in the chirality-induced magnetization with increasing temperatures.
MOKE measurements
MOKE microscopy is an optical technique used to measure the surface magnetization of a material by reflecting polarized light from its surface, producing image contrast that is directly proportional to the relative magnetization [kapon2023magneto, ozturk2023magnetization, sharma2020control]. MOKE microscopy was used to measure the chirality-induced avalanche magnetization on a ferromagnetic surface by enantiopure RAO crystals.
Enantiopure RAO crystals were formed on a Si/Ni(50 nm)/Au(5 nm) thin film by drop-casting an RAO water solution, as illustrated in Figure 2a. A commercial Evico-Magnetics GmbH Magneto-optical Kerr effect (MOKE) microscope, equipped with an electromagnet oriented parallel to the sample surface, was used to image the local magnetic effects induced by the enantiopure RAO crystals. The MOKE microscope provides qualitative imaging of surface magnetization by detecting the Kerr rotation, which reflects the change in the linear polarization of light as a function of the local magnetization (Figure 2a). The contrast in the MOKE images corresponds to regions with different magnetization directions relative to the incident light, allowing us to assess the influence of the enantiopure RAO crystals on adjacent magnetic domains.
Surface magnetization images of the Ni film around the enantiopure RAO crystals at different magnetic fields and temperatures (20°C blue, 40°C light blue, 60°C light red, 80°C red) were taken. Magnetic hysteresis loops were obtained from surface magnetization images by sweeping the magnetic field and averaging the intensity in areas far (Figure 2b top) and under (Figure 2b bottom) from the enantiopure RAO crystals within the same image. As the temperature increases, the magnetic hysteresis loops under the crystals widen, indicating an increase in the coercive field. In contrast, the coercive field of the bare magnetic surface far from the crystals remains constant across all temperatures. During the measurements, RAO crystals remain stable at all temperatures.
Starting with a negative field that saturates the sample, a uniform dark magnetization is observed, corresponding to -1 magnetization. As the magnetic field increases, the Ni magnetization flips to bright (+1 magnetization) at 1 mT; however, it requires an additional 2 mT to flip the domains surrounding the RAO crystals. When sweeping the field back down, the domains far from the crystals flip at -1 mT, while those under the crystals flip at -3 mT. The coercive field is plotted as a function of temperature in Figure 2c for regions under (red) and far (gray) from the crystals. The coercive field under the crystal is 2 mT higher than that of the bare surface far from the crystals and increases linearly with temperature.
The optical image of the sample is shown in Figure 2d with the area of the crystals marked in a dashed white line. The areas under the crystals (red) or far from the crystals (gray) from which the magnetization loops were taken are marked on the image. Corresponding magnetic images at a 3.85 mT magnetic field are shown in Figure 2e (20°C) and Figure 2f (80°C). In this field, all the domains flipped at 20°C while at 80°C some light domains are left under the crystals.
At higher temperatures, a more gradual transition from one magnetization to the other under the crystals is seen by the increasingly slow transition toward magnetic saturation in the magnetic hysteresis loops in Figure 2b. This is due to the smaller domain size at higher temperatures. Thus, while stronger magnetization is observed at higher temperatures, the spread of avalanche magnetization decreases, as seen in Figure S6.
Hall effect measurements
When an electric current flows through a conductor in a perpendicular magnetic field, the Lorentz force creates a transverse Hall voltage (). In ferromagnets, the anomalous Hall effect introduces an additional term from intrinsic magnetization. The total Hall voltage is: . Where is the regular Hall term and is the anomalous term due to the material’s magnetization. The anomalous Hall effect is commonly used to probe surface magnetization induced by chiral molecules [kumar2017chirality, eckshtain2016cold].
The magnetization induced by the RAO crystals on a Si/Ni(40 nm)/Au(10 nm) thin film was measured by a simple hall device in Van der Pauw configuration as shown in Figure 3a. Au pads were fabricated at the corners of a Si/Ni(40 nm)/Au(10 nm) thin film, the current was applied (), and the perpendicular Hall voltage () was measured under external magnetic field of -25 mT perpendicular to the surface. Enantiopure RAO was drop-cast onto the substrate and left to dry. The Hall voltage was measured before and after RAO adsorption to avoid any variations between geometry or magnetic properties between samples.
The Hall resistance () was extracted by plotting the Hall voltage as a function of the current and determining the slope from a linear fit to the data. This was calculated at different temperatures, both with and without RAO crystals. In both cases, increases with temperature, but with the RAO crystals it increases much faster. The hall voltage in ferromagnets has contributions from both the external magnetic field and the magnetization. Therefore, at a given magnetic field, the increase in corresponds to increasing magnetization. The slight increase in of the bare substrate was observed before and contributed to the temperature dependence of the intrinsic anomalous Hall effect [Nihallberry, Nihallmag].
The results indicate that the magnetization of the surface increases with temperature in the presence of chiral RAO crystals, consistent with what was observed in the MOKE results. In Figure 3b, of the bare substrate, at 30°C, is while of the substrate with the RAO crystals is . The increase supports the previously established assumption that the chiral molecules themselves are magnetizing the surface.
The role of temperature in RAO synthesis
The spin-controlled processes responsible for the chirality-induced magnetization by RAO become significantly more effective at higher temperatures.
Additionally, we studied the influence of temperature on the prebiotic synthesis of RAO within a mixture of pentose-aminooxazolines, isomers of RAO. NMR measurements revealed that higher temperatures led to increased relative yields of RAO, due to its greater thermodynamic stability.
NMR measurements
A 1:1 mixture of DL-glyceraldehyde and 2-aminooxazole (2-AO), each at a concentration of 250 mM, was prepared in deionized water. The solution was then divided among five tubes, with each tube incubated at a different temperature: 5°C, 20°C, 35°C, 50°C, and 65°C. The samples were analyzed by 1H nuclear magnetite resonance (NMR) after 2.5 hours and again after 24 hours and the absolute concentrations of compounds were determined using a quantitative NMR standard.
The prebiotic synthesis discovered by the Sutherland group produced four stereoisomers—ribose-, arabinose-, xylose-, and lyxose-aminooxazolines—along with their respective enantiomers [powner2009synthesis] (Figure 4a). Of these, ribose-aminooxazoline (RAO), the precursor to RNA, is the least soluble and crystallizes as a conglomerate. The major isomer, RAO, was produced alongside others, with arabinose-aminooxazoline (AAO) being nearly as abundant. When the reactions were completed, the pentose-AOs had an overall yield of 95%. Notably, the ratios of AOs varied across different temperatures.
Temperature influenced the reaction in two ways: it accelerated both the synthesis of pentose-AOs and, once synthesis was complete, their degradation. After 2.5 hours, before the reaction was completed, a slight increase in the RAO to AAO ratio was observed from 1.2 to 1.3 (Figure 4b). After 24 hours, when the reaction was completed for temperatures 20°C above, degradation of pentose-AO by hydrolysis to their corresponding oxazolidinones was observed. Interestingly, this degradation led to an increase in the RAO/AAO ratio from 1.2 to 1.5 from 20°C to 65°C, marking a 23% increase of the desired isomer, RAO, relative to the second most abundant isomer, AAO (Figure 4c). The observed selective degradation favoring RAO over AAO is attributed to RAO’s greater thermodynamic stability—consistent with the earlier study of isomer conversion mediated by phosphate as a catalyst [powner2010phosphate]. Consequently, RAO’s abundance in the pentose-AO mixture increased by 14%, from 44% to 50%. Nevertheless, the RAO-to-XAO ratio decreased from 5.3 to 3.86 (Figure 4c), indicating that selective degradation by heating does not mitigate the unfavorable effects of XAO during crystallization, as observed in our previous study [ozturk2023crystallization]. It is also important to note that the degradation of pentose-AOs by hydrolysis was not facilitated by a general acid-base catalyst, such as phosphate, bicarbonate or bisulfite, which could accelerate the process.
The observation that the relative abundance of the desired isomer, RAO, rises with increasing temperature complements our earlier findings on the improved efficiency of RAO-induced magnetization at higher temperatures. This confirms that our mechanism can effectively operate over a broad temperature range, demonstrating even greater efficiency at higher temperatures.
Discussion
@@
Having provided an intuitive thermodynamic picture in the introduction, we will now present a microscopic model connecting phonons to electron spin, and thereby to magnetization. This model produces the red curve displayed in Figure 1d.
@@
The strength of the coercive field is a measure of the axial magnetic anisotropy, which corresponds to the energy barrier stabilizing a ferromagnet against spin excitations. These spin excitations, or magnons, are essentially destructive to the ferromagnetic order. However, by studying the magnon spectrum, we gain insight into the cause of the temperature-enhanced coercive field observed in the experiments (Figure 2 and Figure 3).
@@
This model is capable of describing the ferromagnetic state of a local moment ferromagnet for temperatures below the Curie temperature, . While the Heisenberg interaction defines a ferromagnetic ground state, this state is spatially isotropic, with the effect that a magnetic moment cannot be derived solely in terms of this interaction. This is also a consequence of the Mermin and Wagner theorem, which states that local isotropic interactions cannot give rise to long-range order for systems with spatial dimensions less than three [mermin1966absence]. The anisotropic contribution relaxes this fundamental principle and introduces an energy gap such that any excited state of the spin system requires additional energy to be occupied—energy that may be provided by the temperature of the surrounding environment. Eventually, this energy gap is closed by the thermal energy at sufficiently high temperatures, leading to the destruction of the ferromagnetic order. This is the reason for the destruction of ferromagnetism with temperature, which is the initial decay seen in red in Figure 3d.
The unexpected increase in magnetization with temperature observed in this work can be explained by the coupling between nuclear vibrations and the localized spin moment [fransson2017microscopic]. This is the origin of the enhanced chirality-induced magnetization with increasing temperatures.
@@
where denotes the spatial displacement of the nucleus at , while represents a coupling between the localized spin moment at and the displacement (or the nucleus at ). The nuclear displacements result in a field acting on the localized spin moments, resembling that of a magnetic field; hence, we may refer to it as a pseudo-magnetic field. It should be noted that is a tensorial interaction with the form of a susceptibility, which arises from charge-spin fluctuations in the compound. While this susceptibility can be considered negligible in inversion-symmetric crystal structures, is non-vanishing in chiral structures [fransson2023chiral], where inversion symmetry is broken by definition. Introducing this component to the model is relevant in the current context, where we have chiral molecules adsorbed on a magnetite surface.
As we are interested in the effect of the spin-lattice coupling on the magnetic anisotropy, we consider the linear spin wave spectrum that results from the introduced model. The details will be presented in a follow-up work, and here, we merely present the salient features from this investigation.
@@
At higher temperatures, the dominant processes that govern the physics are related to the spin-lattice coupling, such that the energy dispersion relation must be replaced by , where the additional anisotropic contribution provides a linear temperature dependence of the energy dispersion, as seen in the shaded blue area of the red curve in Figure 1d. The effect of this linear contribution to the anisotropy energy leads to an increase in the coercive field with temperature, in agreement with the experimental observations (Figure 2 and Figure 3). Physically, we interpret this addition to the anisotropic energy as a new channel that diverts the energy away from exciting spin waves in the system and instead generates lattice vibrations. This term is connected to heat supplied to the chiral engine, as we explained in the thermodynamic picture.
@@
The combination of these two terms—one being destructive and the other constructive to the thermodynamic order—stabilizes the system of chiral molecules on a magnetic surface against environmental fluctuations. This provides persistent magnetization necessary for observing spin-controlled asymmetric processes due to CISS in natural environments. This inherent stability of our mechanism, together with RAO’s superior thermodynamic stability (Figure 4), enhances our scenario’s environmental plausibility.

### v3

# Temperature-Enhanced Coercive Field by Chiral Molecules

@@ Abstract
The chiral-induced spin selectivity (CISS) effect demonstrates a strong coupling between electron spin and molecular chirality, enabling spin-controlled interactions between chiral molecules and magnetic surfaces. While CISS experiments have revealed robust changes in the spin-polarization properties of magnetic materials upon chiral molecular adsorption, the temperature dependence of these effects remains poorly understood. Here, we investigate the temperature dependence of the chirality-induced increase in magnetic coercivity by ribose-aminooxazoline (RAO) crystals on ferromagnetic surfaces. RAO was selected as a conglomerate-forming, thermodynamically stable crystalline chiral organic molecule with prebiotic relevance that has previously been shown to induce strong spin-dependent changes in magnetic minerals. Contrary to classical expectations that magnetic coercivity weakens at elevated temperatures, we observe a significant increase in magnetic coercivity (1mT over a 60∘C temperature change) with increasing temperature. These results support a vibronic contribution to CISS arising from electron–phonon interactions and demonstrate that spin-dependent interactions between chiral molecules and magnetic surfaces can become more effective at higher temperatures, providing new insight into the microscopic origins of CISS and the environmental robustness of spin-controlled asymmetric processes.
Keywords: chirality, magnetism, CISS effect, temperature
@@
The chiral-induced spin selectivity (CISS) effect—is an empirical phenomenon revealing a strong interaction between electron spin and molecular chirality, even at room temperature [25, 24, 6]. The strong spin-chirality coupling due to CISS [40, 33] allows for achiral magnetic minerals to function as chiral reagents for chemical reactions, which can be controlled by electron spin [3, 5, 22]. Concurrently, the reciprocal nature of this effect was demonstrated by showing that chiral molecules can induce magnetization on magnetic thin films [27, 17, 4] and locally change the magnetic anisotropy of thin films [36, 23](Figure 1a). Although various independent measurements confirm CISS-related effects, a complete microscopic understanding of CISS remains an ongoing challenge [9]. Notably, recent theoretical work by Fransson emphasizes the role of electron correlations in explaining the high levels of spin-polarization observed in experiments [12].
@@
Additionally, the effect of temperature on CISS through electron-phonon interactions presents an intriguing complexity [6, 18, 13], with experimental results showing trends that appear incompatible with each other [7, 32, 2, 39]. This complexity is partly due to the fact that temperature also affects the material and molecular properties involved, complicating efforts to isolate the CISS component of the observed effects. To address these challenges, we examined the temperature dependence of chirality-induced avalanche magnetization in chiral ribose-aminooxazoline (RAO) crystals on magnetic surfaces stable within our temperature range (Figure 1). RAO is a conglomerate-forming central RNA and DNA precursor of prebiotic relevance, characterized by high thermodynamic stability and a well-established crystal structure, and it exhibits strong spin-dependent interactions with magnetic surfaces [35, 29, 30, 38, 28].
@@
In contrast to classical expectations for spin-controlled processes, our experimental results and theoretical analysis show that the adsorption of chiral molecules leads to a significant increase in magnetic coercivity—via their spin-selective interactions with the magnetic surface—that becomes more enhanced at higher temperatures (Figure 2). This supports a vibronic contribution to CISS, indicating that phonon-assisted couplings play a significant role in enhancing the ferromagnetic order due to the CISS effect [10]. These findings shed light on the temperature dependence of CISS, which remains an active area of research with seemingly incompatible findings.
Results
We studied the effect of temperature on chirality-induced magnetization using a magneto-optical Kerr effect (MOKE) microscope and showed an increase in magnetic coercivity with increasing temperature.
MOKE microscopy is an optical technique used to measure the surface magnetization of a material by reflecting polarized light from its surface, producing image contrast that is directly proportional to the relative magnetization [17, 27, 34]. MOKE microscopy was used to measure the chirality-induced avalanche magnetization on a ferromagnetic surface by enantiopure RAO crystals.
Enantiopure RAO crystals were formed on a Si/Ni(50 nm)/Au(5 nm) thin film by drop-casting an RAO water solution, as illustrated in Figure 2a. A commercial Evico-Magnetics GmbH Magneto-optical Kerr effect (MOKE) microscope, equipped with an electromagnet oriented parallel to the sample surface, was used to image the local magnetic effects induced by the enantiopure RAO crystals. The MOKE microscope provides qualitative imaging of surface magnetization by detecting the Kerr rotation, which reflects the change in the linear polarization of light as a function of the local magnetization (Figure 2a). The contrast in the MOKE images corresponds to regions with different magnetization directions relative to the incident light, allowing us to assess the influence of the enantiopure RAO crystals on adjacent magnetic domains. It is important to note that MOKE microscopy here is not used for measuring induced magnetization upon adsorption of RAO molecules, but the coercive field of the sample with and without the RAO molecules, i.e., the sample’s response to an outer magnetic field.
Surface magnetization images of the Ni film around the enantiopure RAO crystals at different magnetic fields and temperatures (20°C blue, 40°C light blue, 60°C light red, 80°C red) were taken. Magnetic hysteresis loops were obtained from surface magnetization images by sweeping the magnetic field and averaging the intensity in areas far (Figure 2b top) and under (Figure 2b bottom) from the enantiopure RAO crystals within the same image. As the temperature increases, the magnetic hysteresis loops under the crystals widen, indicating an increase in the coercive field. In contrast, the coercive field of the bare magnetic surface far from the crystals remains constant across all temperatures. During the measurements, RAO crystals remain stable at all temperatures. Racemic RAO was not studied since RAO crystallizes as a conglomerate, inevitably creating homochiral domains. Achiral crystals of glycine and sodium chloride were studied under the same conditions as control experiments (Figure 3). Unlike the chiral samples, they showed no change in coercivity relative to the bare substrate, further confirming that the observed effect arises from molecular chirality.
Starting with a negative field that saturates the sample, a uniform dark magnetization is observed, corresponding to -1 magnetization. As the magnetic field increases, the Ni magnetization flips to bright (+1 magnetization) at 1 mT; however, it requires an additional 2 mT to flip the domains surrounding the RAO crystals. When sweeping the field back down, the domains far from the crystals flip at -1 mT, while those under the crystals flip at -3 mT. A video of the MOKE images during a single sweep is presented in Videos S1 and S2, and an example of the resulting images is presented in Figure S3. The coercive field is plotted as a function of temperature in Figure 2c for regions under (red) and far (gray) from the crystals. The coercive field under the crystal is 2 mT higher than that of the bare surface far from the crystals and increases linearly with temperature.
The optical image of the sample is shown in Figure 2d with the area of the crystals marked in a dashed white line. The areas under the crystals (red) or far from the crystals (gray) from which the magnetization loops were taken are marked on the image. Corresponding magnetic images at a 3.85 mT magnetic field are shown in Figure 2e (20°C) and Figure 2f (80°C). In this field, all the domains flipped at 20°C, while at 80°C, some light domains are left under the crystals. Both enantiomers of RAO were studied, and they led to similar phenomena of increased magnetic coercivity, as presented in Supporting Information Section 1.2.
At higher temperatures, a more gradual transition from one magnetization to the other under the crystals is seen by the increasingly slow transition toward magnetic saturation in the magnetic hysteresis loops in Figure 2b. This behavior likely arises from reduced domain size at higher temperatures. Consequently, although the coercive field is stronger, the extent of avalanche magnetization decreases (Supporting Information Section 1.5), highlighting a tradeoff between the strength of induced magnetization and its spatial spread with increasing temperatures.
Discussion
@@
How can we account for the increased effectiveness of CISS with temperature without involving microscopic details? We will start with an intuitive thermodynamic picture and incorporate microscopic analyses later in the discussion.
CISS relates to the spin-filtering of electrons, meaning that chiral molecules reduce entropy by decreasing the number of accessible microstates of the electronic spin distribution at the molecule–surface interface.
We first consider the electronic spin distribution as the system of interest. To leading order, electron number and energy are conserved during spin filtering, allowing this subsystem to be described within a microcanonical ensemble. According to the first law of thermodynamics,
Spin filtering reduces the number of accessible spin microstates, implying for the electronic subsystem. This entropy reduction requires positive work to be performed by the chiral molecules.
We next consider the chiral molecules as a separate subsystem interacting with their surroundings. Since the entropy of the electronic subsystem decreases, the entropy of the chiral molecules must increase according to the second law of thermodynamics. The free energy change of the molecular subsystem is
Because for the molecules, the term becomes increasingly negative at higher temperatures. Since the spin filtering process is not spontaneous, heat must be supplied to the molecular subsystem to maintain a positive free energy cost.
In this picture, chiral molecules operate as Carnot-like engines, absorbing heat from their surroundings and converting it into work that reduces the entropy of the electronic spin distribution (Figure 4a). Consequently, these chiral engines operate more effectively when heat is provided.
@@
This thermodynamic picture is an intuitive way of visualizing the temperature dependence of CISS. However, the effect itself is a complex phenomenon involving spin-orbit coupling, electron correlations, and electron spin-phonon interactions. Fransson investigated the role of phonons and vibronic interactions in CISS, constructing a microscopic theory of CISS beyond the Born–Oppenheimer approximation. This model elucidates the exchange of angular momentum as electrons traverse a helical trajectory [13, 14], which is introduced below and elaborated upon in our complementary theoretical study [10]. Our experimental results align with this microscopic framework, indicating that phonon-chirality couplings, amplified by temperature, play an essential role in CISS.
Microscopic model
@@
Having provided an intuitive thermodynamic picture, we will now present a microscopic model connecting phonons to electron spin, and thereby to magnetization. This model produces the curves displayed in Figure 4b.
@@
The strength of the coercive field is a measure of the axial magnetic anisotropy, which corresponds to the energy barrier stabilizing a ferromagnet against spin excitations. These spin excitations, or magnons, are essentially destructive to the ferromagnetic order. However, by studying the magnon spectrum, we gain insight into the cause of the temperature-enhanced coercive field observed in the experiments (Figure 2).
@@
This model is capable of describing the ferromagnetic state of a local moment ferromagnet for temperatures below the Curie temperature, . While the Heisenberg interaction defines a ferromagnetic ground state, this state is spatially isotropic, with the effect that a magnetic moment cannot be derived solely in terms of this interaction. This is also a consequence of the Mermin and Wagner theorem, which states that local isotropic interactions cannot give rise to long-range order for systems with spatial dimensions less than three [21]. The anisotropic contribution relaxes this fundamental principle and introduces an energy gap such that any excited state of the spin system requires additional energy to be occupied—energy that may be provided by the temperature of the surrounding environment. Eventually, this energy gap is closed by the thermal energy at sufficiently high temperatures, leading to the destruction of the ferromagnetic order. This is the reason for the destruction of ferromagnetism with temperature.
The unexpected increase in magnetic coercivity with temperature observed in this work can be explained by the coupling between nuclear vibrations and the localized spin moment [11]. This is the origin of the enhanced chirality-induced magnetization with increasing temperatures.
@@
where denotes the spatial displacement of the nucleus at , while represents a coupling between the localized spin moment at and the displacement (or the nucleus at ). The nuclear displacements result in a field acting on the localized spin moments, resembling that of a magnetic field; hence, we may refer to it as a pseudo-magnetic field. It should be noted that is a tensorial interaction with the form of a susceptibility, which arises from charge-spin fluctuations in the compound. While this susceptibility can be considered negligible in inversion-symmetric crystal structures, is non-vanishing in chiral structures [15], where inversion symmetry is broken by definition. Introducing this component to the model is relevant in the current context, where we have chiral molecules adsorbed on a magnetite surface.
As we are interested in the effect of the spin-lattice coupling on the magnetic anisotropy, we consider the linear spin wave spectrum that results from the introduced model. The details are presented in a complementary work [10], and here, we merely present the salient features from this investigation.
@@
At higher temperatures, the dominant processes that govern the physics are related to the spin-lattice coupling, such that the energy dispersion relation must be replaced by , where the additional anisotropic contribution provides a linear temperature dependence of the energy dispersion, as seen in the plotted curve in Figure 4b. The derivation of this linearity is discussed in [10] and in the SI in section 4. The effect of this linear contribution to the anisotropy energy leads to an increase in the coercive field with temperature, in agreement with the experimental observations (Figure 2). Physically, we interpret this addition to the anisotropic energy as a new channel that diverts the energy away from exciting spin waves in the system and instead generates lattice vibrations. This term is connected to heat supplied to the chiral engine, as we explained in the thermodynamic picture.
@@
The combination of these two terms—one being destructive and the other constructive to the thermodynamic order—stabilizes the system of chiral molecules on a magnetic surface against thermal fluctuations. This provides persistent magnetization necessary for observing spin-controlled asymmetric processes due to CISS in natural environments. This inherent stability of our mechanism, together with RAO’s superior thermodynamic stability [31], enhances our scenario’s environmental plausibility.
Increased coercivity and the role of magnetic substrate
The magnetic response in our system manifests primarily as an increase in coercive field, rather than as a reversal of the net magnetization. This behavior differs from earlier reports of magnetization switching in perpendicular-anisotropy stacks [4], which are distinct from both our in-plane Ni/Au films and bulk magnetite.
The MOKE loops remain symmetric in our measurements and show no evidence of exchange bias, though we cannot fully exclude subtle effects below our detection threshold. This observation supports a scenario of anisotropy change rather than molecule-driven magnetization switching. A possible explanation is that the molecules induce local magnetic moments at the surface [1]. Such surface effects can alter magnetic anisotropy but may not be sufficient to reverse the magnetization of the ferromagnet within an exchange length.
To distinguish between coercivity enhancement and true magnetization switching, we performed complementary Hall effect measurements on Ni films with Au or oxide capping layers (Supporting Information Section 2). These measurements show an increase in Hall resistance upon RAO adsorption; however, the response lacks consistent directionality, likely due to nonspecific molecular binding at the surface. Therefore, based on our current measurements, we cannot conclude that RAO induces net magnetization switching in Ni/Au systems. However, our theoretical model allows us to probe the net magnetization induced by chiral molecules and its temperature dependence. It predicts that , the net out-of-plane magnetization, decreases with increasing temperature [10], unlike the magnetic coercivity, which is enhanced with temperature in agreement with our measurements.
Surface chemistry appears central to these considerations. On magnetite, the magnetic surface that was studied in connection to prebiotic chemistry of RAO, RAO directly binds through hydrogen bonding, a strong chemical interaction that brings the molecules into close proximity with the substrate. In contrast, on 5 nm gold-capped nickel, RAO interacts more weakly and less directly through electrostatic forces. This weaker physisorption also keeps the molecules farther from the surface, leading to reduced interfacial coupling and less organized dipole orientation. Molecular orientation is particularly important: as shown by Nguyen et al. [26], oppositely oriented molecules can invert the direction of the CISS effect, and strong collective co-alignment of molecules is required for a strong CISS signal.
Nickel surfaces introduce an additional complication. Their metallic nature induces a mirror potential that screens surface charges, effectively reducing the influence of adsorbed molecules on spin-dependent charge redistribution due to CISS [20]. This screening effect should be weaker on semiconducting magnetite, allowing stronger proximity interactions with chiral molecules.
In summary, RAO adsorption on magnetite produces a more pronounced CISS effect than on Ni/Au. This enhancement arises from (i) stronger chemical binding, (ii) stronger alignment of dipoles, (iii) closer molecular proximity to the surface, and (iv) weaker electronic screening by the semiconducting substrate.
Our earlier studies of RAO on magnetite, which revealed a handedness-dependent inversion of the circular dichroism signal, suggest that the RAO–magnetite interface can produce a net magnetization change within the magnetite exchange length (around 9 nm). In contrast, MOKE measurements of RAO on Ni/Au show no evidence of exchange bias in the hysteresis loops [27]. Taken together, these results clarify the subtle differences in the magnetic behavior of RAO on magnetite versus Ni/Au and provide a mechanistic basis for the enhanced CISS effect observed on magnetite.
Last but not least, it is also important to consider the role of magnetic domain size and structure. Our MOKE measurements demonstrate that the observed effects depend on magnetic domain size. In multidomain thin films, such as those studied here, in-plane magnetization switching requires a strong induced effective magnetic field, while coercivity can increase more readily through local changes in anisotropy. Under these conditions, any induced magnetization is transient and disappears once the molecules are removed. In contrast, systems with frustrated geometries—such as perpendicular-anisotropy stacks or single-domain particles—are more susceptible to stable magnetization reversal. Single-domain, low-coercivity magnetite particles represent the most suitable magnetic substrates for our system. It is therefore important to investigate whether RAO can induce permanent magnetization switching in these low-coercivity grains. These particles are also geochemically plausible under early Earth and Martian conditions and can be synthesized prebiotically from iron–carbonate solutions [16, 37].